Solve using the five-step problem-solving process. Show all steps necessary to arrive at your solution.
A rectangular board is 8 inches longer than it is wide. If the area of the board is 48 square inches, find the length and the width.

Answer by helpnalgebra(91)

x(x+8)=48
x^2+8x=48
x^2+8x-48=0
x^2+12x-4x-48=0
x(x+12)-4(x+12)=0
(x+12)(x-4 ) = 0
x= -12 Or 4
So width = 4 inches
Length =12 inches